OpenStudy (anonymous):

A 14.01 g sample of N2 reacts with 3.02 g of H2 to form ammonia (NH3). If ammonia is the only product, what mass of ammonia is formed?

OpenStudy (cuanchi):

1) write down and balance the equation a H2 + b N2 -> c NH3 2) calculate the moles of N2 in 14.01 g of N2 and the number of moles of H2 you have in 3.02g of H2 14.01 g x 1mol N2/ 14.01 g= moles of N2 3.02 g H2 x 1 mol N2/ 1.01g = moles of H2 3) according to the stoichiometry of the reaction you can calculate how many moles of NH3 you can produce with the moles of N2 and the amount of NH3 that you can produce with the moles of H2. moles of N2 x c/b = moles of NH3 moles of H2 x c/a = moles of NH3 the smallest of both of these values will be the maximum amount of NH3 that you can produce with that amounts of N2 and H2 and also that element is going to be the limiting reactant of the reaction